assignment 1

Make sure you are able to write a python program as described in the text. You will be working in repl.it

  • Title your program temperature.
  • Add a comment to say that this program converts Fahrenheit to Centigrade.
  • Create variables:
    • fahrenheit and centigrade values
    • my name
  • Have the program prompt the user (the instructor) for a value for the fahrenheit variable.
  • Convert the value to centigrade, using this formula:
    (fahrenheit - 32) * ÷ 9 = centigrade
  • Have the program print results (the centigrade value)to the screen. Concatenate the value for the my name variable in to the screen print.

assignment 2

  • Title your program temperature2.
  • Add a comment to say that this program allows conversion to either Fahrenheit or Centigrade.
  • Create variables:
    • fahrenheit and centigrade values
    • my name
    • and any other variables you may need
  • Have the program ask the user (the instructor) if they want to convert to Fahrenheit or Centigrade
  • Based on response, run the appropriate formula for conversion:
    • (fahrenheit - 32) * 5 ÷ 9 = centigrade
    • centigrade * 9 ÷5 + 32 = fahrenheit
Answer #1
# assignment 1
# this program converts Fahrenheit to Centigrade.
fahrenheit = 0
centigrade = 0
my_name = 'ronaldo'

# Have the program prompt the user (the instructor) for a value for the fahrenheit variable.
fahrenheit = float(input("Enter temperature in fahrenheit: "))
# Convert the value to centigrade, using this formula:
# (fahrenheit - 32) * (5 / 9) = centigrade
centigrade = (fahrenheit - 32) * (5 / 9)
# Have the program print results (the centigrade value)to the screen.
# Concatenate the value for the my name variable in to the screen print.
print(my_name + ": " + str(fahrenheit) + " fahrenheit is " + str(centigrade) + " centigrade")

# assignment 2
# this program converts either Centigrade to Fahrenheit or Fahrenheit to Centigrade.
fahrenheit = 0
centigrade = 0
my_name = 'ronaldo'

# Have the program ask the user (the instructor) if they want to convert to Fahrenheit or Centigrade
print("1. Fahrenheit to Centigrade")
print("2. Centigrade to Fahrenheit")
choice = int(input("Enter your choice(1 or 2): "))
if choice == 1:
    fahrenheit = float(input("Enter temperature in fahrenheit: "))
    # Convert the value to centigrade, using this formula:
    # (fahrenheit - 32) * (5 / 9) = centigrade
    centigrade = (fahrenheit - 32) * (5 / 9)
    # Have the program print results (the centigrade value)to the screen.
    # Concatenate the value for the my name variable in to the screen print.
    print(my_name + ": " + str(fahrenheit) + " fahrenheit is " + str(centigrade) + " centigrade")
else:
    centigrade = float(input("Enter temperature in centigrade: "))
    # Convert the value to centigrade, using this formula:
    # (fahrenheit - 32) * (5 / 9) = centigrade
    fahrenheit = centigrade * 1.8 + 32
    # Have the program print results (the centigrade value)to the screen.
    # Concatenate the value for the my name variable in to the screen print.
    print(my_name + ": " + str(centigrade) + " centigrade is " + str(fahrenheit) + " fahrenheit")